I really need help on this question:
"Write the equation of the line (in slope intercept form) that goes through the points (7,-4) and has slope=1"
Also
"Write the equation of the line (in slope intercept form) that goes through the points (2,-7) and (2,5)"
Thank you so much!!!!

Answer by Cromlix(4381)

Using equation: y - b = m(x - a)
m = 1 (7, -4)
y -(-4) = 1(x - 7)
y + 4 = x - 7
y = x - 7 - 4
y = x - 11
...............
Using gradient equation to find slope (m)
y2 - y1/x2 - x1
5 - (-7)/2 - 2
12/0
As the denominator = 0
then this shows that the line is a straight line
x = 2.
Hope this helps.
:-)